# Antipersistent energy current correlations in strong long-ranged Fermi-Pasta-Ulam-Tsingou type models

### 17 Jan 2024 v1

@@ V Scattering dynamics of phonons and DBs
The subdiffusive transport mechanism in the mean-field case was understood by DBs LRHeat-2017-2 ; LRHeat-ours-2 . Therefore, it is also interesting to study phonon-DBs scattering in the systems with LR interactions beyond the mean-field regime. From the above figures in Secs. III and IV, it seems that the second law of $\rho_{Q}(0,t)\sim t^{-\gamma}$ emerges at around $\sigma\approx 0.8$. Thus, such an examination of dynamics may help us understand the origin and provide further insight.
To capture the phonon-DBs scattering dynamics in our systems, we perform such a numerical experiment: At time $t=0$ when all particles are at their equilibrium positions (the system is not thermalized), we first apply two kicks at the locations of $i=1$ and $i=2205$ with momenta $p_{1}=-0.7$ and $p_{2205}=0.7$ (corresponding to an averaged temperature close to $T=0.5$), respectively. We then carefully examine the evolution of the dynamics of the system. A vivid animation of this phonon-DBs scattering dynamics for several $\sigma$ for a time lag of $\delta t=10$ showing all particle’s momentum $p_{i}$ versus $i$ is given in the Supplementary Material (SM) SM . The relevant snapshots at a long time $t=7000$ is plotted in Fig. 4. As can be seen in both SM and Fig. 4, these phonon-DBs scattering dynamics suggest that the excitations of these systems can be phonons and two kinds of DBs (the standing and moving ones), however, $\sigma\approx 0.7$ is a demarcation point, where the standing DBs are converted to the moving ones [see Fig. 4(f)]. Since the moving DBs can enhance thermal transport LRHeatreview-3 ; LRHeat-ours-1 , this turning point naturally answers the origin of the emergence of the second decay law of $\rho_{Q}(m,t)$. The slight deviation between $\sigma\approx 0.8$ and $\sigma\approx 0.7$ may be caused by that we considered a finite-temperature system for deriving energy current and heat correlations, but we deal with a zero-temperature system to depict phonon-DBs scattering dynamics. In such setups, a kick with momentum of $|0.7|$ might not really correspond to an averaged temperature of $T=0.5$. In addition to this main information, we would also like to note that from the movie in SM SM for $\sigma_{c}=0.5$, if compared to other $\sigma$ values, one can see a standing DBs with the lowest frequency. This additional property is also a sign of the perculiarity of $\sigma_{c}=0.5$, and is consistent with our observations in Secs. III and IV.
@@ VI Conclusion
To summarize, we have studied the thermal transport properties of the Fermi-Pasta-Ulam-Tsingou (FPUT) model under long-range (LR) interactions, particularly within the challenging strong LR regime (where $0\leq\sigma\leq 1$) that eludes current analytical methods. Our findings uncover the intriguing properties of thermal transport in this regime. Specifically, we observed that the energy current correlation is predominantly antipersistent and varies significantly with $\sigma$, exhibiting nonmonotonic behavior. A critical point emerges at $\sigma_{c}=1/2$, where this correlation reaches its minimum negative value. Correspondingly, the heat spread is also dependent on $\sigma$, displaying two distinct scaling laws separated by $\sigma_{c}=1/2$. Intriguingly, these phenomena appear to be linked to the unique $\sigma$-dependent dynamics of phonon-dislocation breather (DB) scatterings. These insights offer a fresh perspective on thermal transport in LR interacting systems, extending beyond mean-field and weak LR regimes.
@@ Acknowledgements.
D.X. is supported by NNSF (Grant No. 12275116) of China, NSF (Grant No. 2021J02051) of Fujian Province of China, and the start-up fund of Minjiang University; J.W. is supported by NNSF (Grant No. 12105122) of China.

### 22 Mar 2024 v3

@@ V Absence of antipersistent energy current correlation
We finally try to give a qualitative explanation for why the absence of antipersistent energy current correlation occurs. In order to achieve this, we examine the phonon-DBs scattering in a zero-temperature system within the whole range of $0\leq\sigma\leq 1$. This scattering dynamics have been employed for demonstrating the absence of subdiffusive transport and antipersistent energy current correlation in the mean-field system with truncated long-range interactions, wherein dominant moving excitations emerge when the truncation is sufficiently strong LRHeat-ours-2 . Here, based on our analysis presented in Secs. III and IV, it has become evident that the emergence of the long tail happens at a value of $\sigma>0.5$, which appears to be closely linked to the absence of antipersistent energy current correlation. Therefore, conducting further dynamic examinations may provide valuable insights and enhance our understanding regarding its origin.
To capture the phonon-DBs scattering dynamics in our systems, we perform such a numerical experiment: at time $t=0$ when all particles are at their equilibrium positions (the system is not yet thermalized), we first apply two kicks at the locations of $i=1$ and $i=2205$ with momenta $p_{1}=-0.7$ and $p_{2205}=0.7$, respectively (it may correspond to an averaged temperature close to $T=0.5$). We then carefully observe the evolution of the dynamics of the system. A visually engaging animation illustrating the phonon-DBs scattering dynamics for various $\sigma$ values with a time lag of $\delta t=10$, displaying momentum $p_{i}$ versus $i$ for all particles, is given in the Supplementary Material (SM) SM . The relevant snapshots at a long time $t=7000$ are depicted in Fig. 4. Before going to discussion of the details, we should note that such a numerical experiment is based on certain specific initial conditions of our setup. It cannot ensure that some of the nonlinear excitations like DBs are sensitive to other initial conditions. In fact, it has been suggested that the stability and mobility of DBs in similar numerical experiments would crucially depend on initial conditions DBsMobile-1 ; DBsMobile-2 . Therefore, our following results can only provide a qualitative but still interesting explanation. Indeed, as evident from both SM and Fig. 4, these scattering dynamics suggest that the excitations in these systems can be mainly attributed to phonons and two types of DBs (standing and moving), while for a value of $\sigma>0.5$, probably being $\sigma\simeq 0.7$ [see Fig. 7(f)], we identify that the standing DBs can be converted into moving ones. This conversion naturally and qualitatively explains the origin of the long tail and the absence of antipersistent energy current correlations. It is also consistent to the observations of Fig. 1 and Fig. 6(c), wherein a similar value of $\sigma>0.5$ being $\sigma\simeq 0.8$ has been revealed. As to the slight difference between the values of $\sigma\simeq 0.8$ and $\sigma\simeq 0.7$, it is interesting to note that previously in Fig. 1 and Fig. 6(c), we considered a finite-temperature system, while here we deal with a zero-temperature system. In this respect, applying kicks with momentum $|0.7|$ in the latter may not capture all information about a thermalized system in the former, even though one might imagine that for a thermalized system with $T=0.5$, all particles are subjected to many kicks with an average momentum magnitude $|0.7|$ but exhibiting fluctuations since $\sqrt{0.5}\approx 0.7$. This demonstrates again that the revealed absence of antipersistent energy current correlations and the microscopic phonon-DBs scattering dynamics is only qualitatively correlated. In addition to this central information, we are also interesting to find that the movie in SM SM reveals an additional distinct characteristic of $\sigma_{c}=0.5$ when compared to other values of $\sigma$, i.e., the presence of standing DBs with the lowest frequency. This additional property may further emphasize the peculiarity associated with $\sigma_{c}=0.5$. However, we acknowledge that currently, we are unable to gain more evidence based solely on the scattering dynamics if compared to the more comprehensive observations provided in Sec. IV.
@@ VI Conclusion
To summarize, we have studied the thermal transport properties of the Fermi-Pasta-Ulam-Tsingou (FPUT) model under long-range (LR) interactions, particularly within the challenging strong LR regime (where $0\leq\sigma\leq 1$) that eludes current analytical methods. Our findings uncover the intriguing properties of thermal transport in this regime. Specifically, we observed that the energy current correlation is predominantly antipersistent and varies significantly with $\sigma$, exhibiting nonmonotonic behavior. An important change point emerges at $\sigma_{c}=0.5$, where this correlation reaches its minimum negative value. Correspondingly, the heat spread is also dependent on $\sigma$, displaying certain special scalings, separated by $\sigma_{c}=0.5$. For $\sigma\rightarrow 0$, a two-stage exponential decay of the central peak of heat spread can be observed, while for $\sigma\rightarrow 1$, a power-law decay with a long tail over extended time periods emerges. Interestingly, this long tail seems starting to emerge at a value of $0.5<\sigma\leq 1$, supporting the absence of antipersistent energy current correlations as observed. Intriguingly, all these phenomena appear to be qualitatively linked to the unique dynamics of $\sigma$-dependent phonon-DBs scatterings dynamics. Combining all of these observations, we conjecture $\sigma_{c}=0.5$ to be a change point characterizing whether final surviving energy carriers after the phonon-DBs scattering are DBs or phonons. In this respect, the absence of antipersistent energy current correlations can also be qualitatively related to the observation of standing DBs converting into moving ones at a value of $0.5<\sigma\leq 1$. All these findings provide a fresh perspective on thermal transport in LR interacting systems that extends beyond mean-field and weak LR regimes, even though further detailed examination of the conjecture is still required.
@@ Acknowledgements.
We are grateful to the referees for their valuable comments, which help us improve the presentation. D.X. is supported by NNSF (Grant No. 12275116) of China, NSF (Grant No. 2021J02051) of Fujian Province of China, and the start-up fund of Minjiang University; J.W. is supported by NNSF (Grant No. 12105122) of China.